The formula for calculating the amount of money returned for an initial deposit into a bank account or CD(cetificate of deposit) is given by
a=p(1+r/n)
A is the amount of the return.
P is the principal amount initially deposited.
r is the annual interest rate(expressed as a decimal).
n is the number of compound periods in one year.
t is the number of years.
Carry all calculations to six decimals on each intermediate step, then round the final answer to the nearest cent.
Supposed you deposit $2,000 for 5 years at a rate of 8%.
a)calculate the return (A) if the bank compounds annulally (n=1). Round your answer to the hundredths place.
ANSWER:
SHOW WORK:

B) calculate the return (A) if the bank compounds quarterly (n=4). Round answer to the hundredths place.
ANSWER:
SHOW WORK:

C)Does compounding annually or quarterly yield more interest? Explain why>
ANSWER:QUARTERLY EARNS MORE INTEREST.
EXPLAIN: BECAUSE YOU MAKE INTEREST ON THE INTEREST DOING IT QUARTERLY.

d)if a bank compounds continuously, then the formula used is A=P square root of e where e is a constant and equals approximately 2.7183. Calculate A with continuous compounding. Round to hundredths place.
ANSWER:
SHOW WORK:

E) a commonly asked question is, how long will it take to double my money? At 8% interest rate and continuous compounding, what is the answer? tound your answer to the hundredths place.
ANSWER:
SHOW WORK:

Answer by checkley77(12844)

a=p(1+r/n)YOU ARE MISSING THE POWER TERM IN HIS EQUATION:
A=P(1+R/N)^T*N
A=2,000(1+.08/1)^5*1
A=2,000(1.08)^5
A=2,000*1.469
A=$2,938.00 ANSWER.
----------------------------------
A=2,000(1+.08/4)^5*4
A=2,000(1+.02)^20
A=2,000(1.02)^20
A=2,000*1.486
A=$2,972.00 ANSWER.
----------------------------------
A=Pe^R*T
A=2,000*2.7183^(.08*5)
A=2,000*2.7183^.4
A=2,000*1.492
A=$2,984.00 ANSWER.